Define degree of ionisation of a weak electrolyte and discuss the factors affecting
it
1
Expert's answer
2020-02-20T07:44:42-0500

Degree of ionization is the ratio of the number of ionized molecules and the number of molecules dissolved in water.It can be expressed as a fraction or in percentage.It is generally denoted by "\alpha" .

For example:

"\ \ \ \" "CH_3COOH(aq) \to CH_3COO^-+H^+"

"t=0\ \ \ \" "\ \ c" "\ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ 0 \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ 0"

"t=t\ \ \ \ \ \ \ \ \ \ \ c(1-\alpha)\ \ \ \ \ \ \ \ \ \ \ \ c\alpha\ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ c\alpha"

where "\alpha" is degree of ionization.

 Factors affecting degree of ionization:

"(a)". More is the dilution,higher is the degree of ionization.

"(b)." Temperature- Generally degree of ionization increases with increase in temperature.

"(c)." More is the amount of solvent,more is the ionization.

"(d)." Degree of ionization decreases in presence of a strong electrolyte having a common ion in accordance with Le-Chatlier principle.

"(e)." Nature of solute